OpenStudy (anonymous):

Can someone help me find the derivative of y=(2sqrx +7)^3/ (x^3-3x^2+1)^7

OpenStudy (anonymous):

i would use logarithmic differentiation for something so annoying

OpenStudy (anonymous):

i don't know what that is. I keep doing it, but i think i am making mistakes some were because part of my answer is wrong

OpenStudy (kropot72):

The explanation here should help you: http://www.math.ucdavis.edu/~kouba/CalcOneDIRECTORY/logdiffdirectory/LogDiff.html

OpenStudy (anonymous):

but i can't use that on a test, thats why i'm doing these type of questions. They are on my test and I can only use the chain, power, constant and quotient rule

OpenStudy (anonymous):

\[y= (2\sqrt{x}+7)^{3}//(x ^{3}-3x ^{2}+1)^{7} \]

OpenStudy (kropot72):

Using the quotient rule and the chain rule gives: \[\frac{\frac{3}{\sqrt{x}}(x ^{3}-3x ^{2}+1)(2\sqrt{x}+7)^{2}-7(2\sqrt{x}+7)^{3}(x ^{3}-3x ^{2}+1)^{6}(3x ^{2}-6x)}{(x ^{3}-3x ^{2}+1)^{14}}\]

OpenStudy (anonymous):

ok, thats what I got for the first step when I did it

OpenStudy (anonymous):

go on

OpenStudy (kropot72):

which can be simplified to: \[\frac{(2\sqrt{x}+7)^{2}[\frac{3}{\sqrt{x}}-7(2\sqrt{x}+7)(x ^{3}-3x ^{2}+1)^{5}(3x ^{2}-6x)]}{(x ^{3}-3x ^{2}+1)^{13}}\]

OpenStudy (anonymous):

I factored out (x^3-3x^2+1)^6

OpenStudy (anonymous):

because in the first step don't you have to multiply the derivative of the top by whats on the bottom, so you will have (x^3-3x^2+1)^7. then you factor (x^3-3x^2+1)^6 out of the whole equation .

OpenStudy (kropot72):

The quotient rule can be written as: \[if\ \ y=\frac{u}{v}\ then\ \ \frac{dy}{dx}=\frac{v\frac{du}{dx}-u\frac{dv}{dx}}{v ^{2}}\] In your question \[u=(2\sqrt{x}+7)^{3}\] \[v=(x ^{3}-3x ^{2}+1)^{7}\] The expression\[(x ^{3}-3x ^{2}+1)^{7}\] does not arise as a result of the process of differentiation.

OpenStudy (anonymous):

the way that I learned it is that f(x)=h(x)/g(x) so the derivative will have the form f'(x)=h'(x)g(x)-g'(x)h(x) divided by g(x)^2

OpenStudy (kropot72):

Another way of expressing the quotient rule is as follows: \[(\frac{f}{g})'=\frac{g.f'-f.g'}{g ^{2}}\]

OpenStudy (anonymous):

yea, thats the way i'm using

OpenStudy (kropot72):

It amounts to the same as the rule expressed interms of u and v.

OpenStudy (anonymous):

ok. but can you use the way that i'm using so I can understand it better. Please. I did this question 3 times and I'm still getting weird answers.

OpenStudy (kropot72):

Can you post your initial step here?

OpenStudy (anonymous):

sure . Do you want the steps that I did, up to were I'm stuck?

OpenStudy (kropot72):

Please post as much as possible.

OpenStudy (anonymous):

\[\frac{(3x ^{-1/2})(2\sqrt{x}+7)^{2}(x ^{3}-3x^{2}+1)^{7}-7(x ^{3}-3x ^{2}+1)^{6}(3x ^{2}-6x)(2\sqrt{x}+7)^{3} }{ (x ^{3}-3x ^{2}+1)^{14} }\]

OpenStudy (anonymous):

\[\frac{(x ^{3}-3x ^{2}+1)^{6}(2\sqrt{x}+7)^{2}[(3x ^{-1/2})(x ^{3}-3x ^{2}+1)-7(3x ^{2}-6x)(2\sqrt{x}+7)]}{ (x ^{3}-3x ^{2}+1)^{14} }\]

OpenStudy (anonymous):

\[\frac{(2\sqrt{x}+7)^{2}[(3x ^{-1/2})(x ^{3}-3x ^{2}+1)-7(3x ^{2}-6x)(2\sqrt{x}+7)] }{ (x ^{3}-3x ^{2}+1)^{8} }\]

OpenStudy (anonymous):

thats what I have so far

OpenStudy (kropot72):

All your working is correct. I made a typo in my initial differentiation and carried the error through. Sorry for that :(

OpenStudy (anonymous):

its ok. But the problem is that I don't know what to do after my third step.

OpenStudy (kropot72):

It appears you have an answer to the question and are trying to simplify your result obtained so far to match the answer. Is my understanding of your situation correct?

OpenStudy (anonymous):

yes

OpenStudy (kropot72):

Can you please post the answer so I can see what the target solution is?

OpenStudy (anonymous):

\[-\frac{ 3(2\sqrt{x}+7)^{2}(13x ^{3}-25^{2}-1+49x^{5/2}-98x ^{3/2}) }{ (x ^{3}-3x ^{2}+1)^{8}\sqrt{x} }\]

OpenStudy (kropot72):

Looking at the numerator of what you have so far: \[(2\sqrt{x}+7)^{2}[(3x ^{-\frac{1}{2}})(x ^{3}-3x ^{2}+1)-7(3x ^{2}-6x)(2\sqrt{x}+7)]\] \[(3x ^{2}-6x)=3x(x-2)\] Also the following expression can be taken out as a factor of the two terms inside the square brackets giving: \[(2\sqrt{x}+7)^{2}(3x ^{-\frac{1}{2}})[(x ^{3}-3x ^{2}+1)-(\frac{7}{3x ^{-\frac{1}{2}}})(3x)(2\sqrt{x}+7)(x-2)]\]

OpenStudy (kropot72):

This can be further simplified. Please wait.

OpenStudy (kropot72):

Continuing with the numerator: When the second expression in the square brackets is multiplied out we get: \[14x ^{3}+49x ^{\frac{5}{2}}-28x ^{2}-98x ^{\frac{3}{2}}\] So the whole expression inside the square brackets becomes: \[-13x ^{3}+25x ^{2}+1-49x ^{\frac{5}{2}}+98x ^{\frac{3}{2}}\] If you take out -1 as a common factor of the above expression and include the other terms you will reach the given answer.

OpenStudy (anonymous):

so my answer was right. Could i just leave it like my answer and not do what you did

OpenStudy (kropot72):

Its difficult to say for sure. But in my opinion most people marking would give aound 8.5 out of 10 for your doing as much as you have done so far and 10 out of 10 for the extra effort going into the simplification.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

thank you so much

OpenStudy (kropot72):

You're welcome :)